OpenStudy (anonymous):

Solve the system of equations.2x – 4y = 12 3x – 6y = 21

OpenStudy (anonymous):

2x – 4y = 12.....A 3x – 6y = 21.....B from equation A find the value of x or y its up to your choice and then place it in equation B in this way you will get the answer of A and B

OpenStudy (anonymous):

sorrry you will get the answer of x and y

OpenStudy (anonymous):

@coolaidd did you get it dear??

OpenStudy (anonymous):

do i multiply by 2

OpenStudy (anonymous):

no you dont hav to multiple it by 2......btw is it an elimination question????

OpenStudy (anonymous):

2x-4y=12 place x on one side and rest of values on other side lie this 2x=12+4y x=12+4y/2 x=12/2+4y/2 x=6+2y .....in this way you get the value of x now put this value in second equationto get the value of y like this 3x-6y=21 3(6+2y)-6y=21 18+8y-6y=21 8y-6y=21-18 2y=3 y=3/2 so in this way you get the value of y as well x=6+2y x=6+2(3/2) x=6+3 x=9

OpenStudy (anonymous):

@coolaidd hope you understand it now =)

OpenStudy (anonymous):

thank you :)